# Uncertainty Quantification for Retrieval-Augmented Reasoning

### v1

@@ 6.5. Action Selection
RQ5 investigates the roll of action set in $\text{R}^{2}$C. Figure 6 illustrates the performance of $\text{R}^{2}$C with different action sets. In the first three bars, only a single action is used, meaning that in each generation, the cut point in the reasoning path is chosen randomly, while the action itself remains fixed. The fourth bar represents our main setup, where both the action and the cut point are selected randomly. For Search-R1 and PopQA, the main setup outperforms all other configurations, whereas in HotpotQA, QP performs slightly better. For SelfAsk, the AV action achieves the best performance, with the main setup ranking second. These results indicate that while our main setup is generally robust, there are still potentials to design action configurations better suited to specific RAR systems.
@@ Appendix A Prompt Details
An important component of the $\text{R}^{2}$C method is the Perturbation Actions discussed in Section 4.2. These actions are implemented by prompting the LLM with specific instructions. Figures 7, 8, and 9 show the prompts used for Query Paraphrasing, Critical Rethinking, and Answer Validation, respectively. Moreover, the Answer Validation action includes a Reasoning Path Summarization component, which is also implemented through prompting an LLM. The prompt for this component is shown in Figure 10. Finally, Figure 11 presents the semantic equivalence prompt, which is used to assess the equality of two responses for majority voting and clustering in the Model Selection process described in Section 5.3, following pervious work (Bakman et al., 2024; Yaldiz et al., 2025a; Li et al., 2025b).
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025a) ReSearch (Chen et al., 2025c) Search-R1 (Jin et al., 2025) Avg.
@@
RrrC (Li et al., 2025b) 52.68 45.25 19.42 41.77 40.25 18.71 44.96 41.58 16.43 45.26 52.13 24.44 56.50 54.70 22.19 38.42
P(true) (Kadavath et al., 2022) 56.26 49.53 18.35 56.53 38.82 15.63 54.47 39.75 15.37 60.42 55.03 29.54 61.11 58.72 24.55 42.27
@@ Appendix C Case Study
Figure 1 illustrates the workflow of the UQ for computing the uncertainty score, which consists of two main steps: Most-likely Generation and Multi-Generations. Based on this workflow, we present two case studies—one successful and one failed, in Figures 12 and 13, respectively. In the successful case, the most-likely response is “Thebarton Oval”, which is incorrect. However, in the Multi-Generations step, only one out of ten generated responses is “Thebarton Oval”. According to Equations (1) and (2), this results in an uncertainty score of 0.9, indicating high uncertainty. In contrast, in the failed case, the most-likely response is “Adrian Lyne”, which is correct. Yet, in the Multi-Generations step, only four out of ten responses are “Adrian Lyne”, again leading to a relatively high uncertainty score. In this sample, we observe that the answer validation action performs correctly and generates a response similar to the most-likely one. However, the query paraphrasing action produces different responses, and two of the three critical rethinking actions also yields a different result. This observation supports the discussion in Section 6.5, which highlights the potential impact of action selection configurations—whether applied per sample, per model, or per dataset.

### v2

@@ 6.5. Action Selection
RQ5 investigates the roll of action set in $\text{R}^{2}$C. Figure 6 illustrates the performance of $\text{R}^{2}$C with different action sets. In the first three bars, only a single action is used, meaning that in each generation, the cut point in the reasoning path is chosen randomly, while the action itself remains fixed. The fourth bar represents our main setup, where both the action and the cut point are selected randomly. For Search-R1 and PopQA, the main setup outperforms all other configurations, whereas in HotpotQA, QP performs slightly better. For SelfAsk, the AV action achieves the best performance, with the main setup ranking second. These results indicate that while our main setup is generally robust, there are still potentials to design action configurations better suited to specific RAR systems. To further improve $\text{R}^{2}$C, future work could focus on learning or adaptively selecting perturbations (e.g., via reinforcement learning), which we expect would enhance both effectiveness and efficiency.
6.6. Generalization Across Backbone LLMs
RQ6 investigates the robustness of $\text{R}^{2}$C to changes in the generation model. Table 4 summarizes the performance of UQ methods under alternative LLM backbones. To ensure diversity, we evaluate one open-source LLM, Gemma2-9b-it, [^6] https://huggingface.co/google/gemma-2-9b-it and one closed-source LLM, GPT-4o-mini. We do not evaluate ReSearch and Search-R1 in this setting, as their fine-tuned models are only available for Qwen family. With Gemma2-9b-it as the backbone, $\text{R}^{2}$C outperforms the second-best method, ReaC, by more than 3%, indicating that the effectiveness of $\text{R}^{2}$C is largely independent of the choice of LLM.
For GPT-4o-mini, $\text{R}^{2}$C outperforms ReaC by more than 9%. The P(true) metric cannot be reported in this setting because token-level probabilities are not accessible. In the case of SelfAsk, we observe a different performance ordering, in which RrrC emerges as the second-best method after $\text{R}^{2}$C, most notably on the Musique dataset. This shift is primarily due to the reduced effectiveness of SelfAsk when GPT-4o-mini is used as the backbone compared to Qwen-2.5-7B-Instruct, despite additional prompt engineering. In particular, SelfAsk achieves exact match accuracies of 18.6%, 12.2%, and 2.0% on PopQA, HotpotQA, and Musique, respectively, which are lower than the corresponding results reported in Table 3. Overall, $\text{R}^{2}$C consistently achieves the highest average performance by a substantial margin, demonstrating strong generalization across different LLM backbones.
@@ Appendix A Prompt Details
An important component of the $\text{R}^{2}$C method is the Perturbation Actions discussed in Section 4.2. These actions are implemented by prompting the LLM with specific instructions. Figures 7, 8, and 9 show the prompts used for Query Paraphrasing, Critical Rethinking, and Answer Validation, respectively. Moreover, the Answer Validation action includes a Reasoning Path Summarization component, which is also implemented through prompting an LLM. The prompt for this component is shown in Figure 10. Finally, Figure 11 presents the semantic equivalence prompt, which is used to assess the equality of two responses for majority voting and clustering in the Model Selection process described in Section 5.3, following pervious work (Bakman et al., 2024; Yaldiz et al., 2025a; Li et al., 2025a).
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
@@
RrrC (Li et al., 2025a) 52.68 45.25 19.42 41.77 40.25 18.71 44.96 41.58 16.43 45.26 52.13 24.44 56.50 54.70 22.19 38.42
P(true) (Kadavath and others, 2022) 56.26 49.53 18.35 56.53 38.82 15.63 54.47 39.75 15.37 60.42 55.03 29.54 61.11 58.72 24.55 42.27
@@ Appendix C Case Study
Figure 1 illustrates the workflow of the UQ for computing the uncertainty score, which consists of two main steps: Most-likely Generation and Multi-Generations. Based on this workflow, we present two case studies—one successful and one failed, in Figures 12 and 13, respectively. In the successful case, the most-likely response is “Thebarton Oval”, which is incorrect. However, in the Multi-Generations step, only one out of ten generated responses is “Thebarton Oval”. According to Equations (LABEL:eq:mv) and (1), this results in an uncertainty score of 0.9, indicating high uncertainty. In contrast, in the failed case, the most-likely response is “Adrian Lyne”, which is correct. Yet, in the Multi-Generations step, only four out of ten responses are “Adrian Lyne”, again leading to a relatively high uncertainty score. In this sample, we observe that the answer validation action performs correctly and generates a response similar to the most-likely one. However, the query paraphrasing action produces different responses, and two of the three critical rethinking actions also yields a different result. This observation supports the discussion in Section 6.5, which highlights the potential impact of action selection configurations—whether applied per sample, per model, or per dataset.